IR/THZ DOUBLE RESONANCE SIGNATURES AT ATMOSPHERIC PRESSURE

\underline{DANE J. PHILLIPS}, ELIZABETH A. TANNER, Kratos Defense and Security Solutions - Digital Fusion Solutions Advanced Technologies Division, 5030 Bradford Dr. Building I, Suite 210, Huntsville, AL 35805; HENRY O. EVERITT, Army Aviation and Missile RD&E Center, Weapons Sciences Directorate, Redstone Arsenal, AL 35898; IVAN R. MEDVEDEV, CHRISTOPHER F. NEESE, JENNIFER HOLT, and FRANK C. DE LUCIA, Department of Physics, 191 W. Woodruff Ave., Ohio State University, Columbus, OH 43210 USA.

IR/THz double resonance (DR) spectroscopy, historically used to investigate molecular collision dynamics and THz molecular lasers at low pressures (< 1 Torr), shows promise for trace gas remote sensing at atmospheric pressure. Molecular specificity is obtained through the rare coincidence(s) between molecule-specific ro-vibrational energy levels and CO_{2} laser lines. The resulting molecule-specific, DR-induced, THz spectroscopic signatures strongly depend on the type of rovibrational transition involved (P, Q, or R), the type of vibrational level excited (stretching or bending), and the molecular mass. To illustrate these sensitivities, calculated DR spectra of prototypical molecules such as methyl fluoride, methyl chloride, and methyl cyanide will be discussed. Although atmospheric pressure broadening obfuscates pure rotational spectra, we show how it can enhance the DR signature in two ways: by relaxing the pump coincidence requirement and by adding the DR signatures of multiple nearby transitions. We will present estimates of this enhancement, including cases where the coincidences that produce the strongest DR signatures at atmospheric pressure do not exist at low pressures.